MICROWAVE SPECTRA OF cis-1,3,5-HEXATRIENE AND ITS ^{13}C ISOTOPOMERS; AN r_{s} SUBSTITU-TION STRUCTURE FOR THE CARBON BACKBONE

RICHARD D. SUENRAM, BROOKS H. PATE, ALBERTO LESSARI, JUSTIN L. NEILL, and STEVEN T. SHIPMAN, Department of Chemistry, University of Virginia, Charlottesville, VA 22904; \underline{NORMAN C. CRAIG}, ROBIN A. HOLMES, and MATTHEW C. LEYDEN, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, OH 44074.

Based on the presumed effects of π-electron delocalization, the adjustments in “C=C” and “C-C” bond lengths should increase with increasing length of polyene molecules. Thus, the adjustments in these bond lengths relative to localized bond lengths should increase from butadiene to hexatriene. Despite its weak dipole moment (0.05 D estimate from B3LYP/cc-pVTZ model), microwave spectra have been recorded for cis-1,3,5-hexatriene on the FT pulsed-jet broadband and cavity spectrometers at the University of Virginia. The small inertial defect (Δ = 0.16700 uÅ^{2}) shows that this molecule is planar. Spectra have also been recorded for the three ^{13}C isotopomers in natural abundance on the cavity instrument. An r_{s} substitution structure of the carbon backbone fitted to ground state rotational constants gives 1.343 Å for the C_{1}=C_{2} bond, 1.457 Å for the C_{2}-C_{3} bond, and 1.335 Å for the C_{3}=C_{4} bond in partial agreement with expectations. After synthesis of deuterium isotopomers of hexatriene and investigation of their microwave spectra, we intend to obtain a semi-experimental equilibrium (r_{e}) structure for direct comparison with theoretical predictions and full assessment of the structural effects of π-electron delocalization.